### v1

# UNIFY: Unified Index for Range Filtered Approximate Nearest Neighbors Search

Abstract.
This paper presents an efficient and scalable framework for Range Filtered Approximate Nearest Neighbors Search (RF-ANNS) over high-dimensional vectors associated with attribute values. Given a query vector $q$ and a range $[l,h]$, RF-ANNS aims to find the approximate $k$ nearest neighbors of $q$ among data whose attribute values fall within $[l,h]$. Existing methods including pre-, post-, and hybrid filtering strategies that perform attribute range filtering before, after, or during the ANNS process, all suffer from significant performance degradation when query ranges shift. Though building dedicated indexes for each strategy and selecting the best one based on the query range can address this problem, it leads to index consistency and maintenance issues.
Our framework, called UNIFY, constructs a unified Proximity Graph-based (PG-based) index that seamlessly supports all three strategies. In UNIFY, we introduce SIG, a novel Segmented Inclusive Graph, which segments the dataset by attribute values. It ensures the PG of objects from any segment combinations is a sub-graph of SIG, thereby enabling efficient hybrid filtering by reconstructing and searching a PG from relevant segments. Moreover, we present Hierarchical Segmented Inclusive Graph (HSIG), a variant of SIG which incorporates a hierarchical structure inspired by HNSW to achieve logarithmic hybrid filtering complexity. We also implement pre- and post-filtering for HSIG by fusing skip list connections and compressed HNSW edges into the hierarchical graph. Experimental results show that UNIFY delivers state-of-the-art RF-ANNS performance across small, mid, and large query ranges.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/sjtu-dbgroup/UNIFY.
1. Introduction
In recent years, Approximate Nearest Neighbors Search (ANNS) has drawn great attention for its fundamental role in data mining (Valkanas et al., 2017; Mo et al., 2022), recommendation systems (Suchal and Návrat, 2010), and retrieval-augmented generation (RAG) (Lewis et al., 2020), etc. Numerous ANNS methods (Arora et al., 2018; Muja and Lowe, 2014; Zheng et al., 2020; Fu et al., 2019; Malkov and Yashunin, 2018; Wang et al., 2021a; Tian et al., 2023) have been developed to efficiently retrieve similar unstructured objects (e.g., text, images, and videos) by indexing and searching their high-dimensional feature vectors (Mo et al., 2022). However, ANNS fails to support many real-world scenarios where users need to filter objects not only by feature similarities but also by certain constraints. For example, Google Image Search allows users to upload an image and search for similar images within a specific period. Likewise, e-commerce platforms such as Amazon enable customers to find visually similar products within a price range.
The above queries can be formulated as the range filtered approximate nearest neighbors search (RF-ANNS) queries. Consider a vector dataset where each vector is associated with a numeric attribute (e.g., date, price, or quantity). Given a query vector $q$ and a range $[l,h]$, RF-ANNS returns $q$’s approximate $k$ nearest neighbors among the data whose attributes are within the range $[l,h]$. Several studies (Wei et al., 2020; Wang et al., 2021b, 2023; Yang et al., 2020; Qin et al., 2020) have been carried out on the RF-ANNS problem, which can be categorized into the following three strategies based on when the attribute filtering is performed.
Strategy A: Pre-Filtering. This strategy performs ANNS after the attribute filtering. For example, Alibaba ADBV (Wei et al., 2020) integrates the pre-filtering strategy using a B-tree to filter attributes, followed by a linear scan on the raw vectors or PQ (Jegou et al., 2010) codes. Milvus (Wang et al., 2021b) partitions data by attributes and builds ANNS indexes for subsets. RF-ANNS is done by first filtering out partitions covering the query range, then performing ANNS on subset indexes. This strategy is efficient for small query ranges, but it does not scale well with larger ranges due to the linearly increasing overhead for scanning qualified vectors or indexes.
Strategy B: Post-Filtering. This strategy uses an ANNS index to find $k^{\prime}$ candidate vectors ($k^{\prime}>k$), then filters by attributes to obtain the final top-$k$ results. Vearch (Li et al., 2018) and NGT (Yahoo, 2016) apply this strategy, which can be easily extended to popular ANNS indexes like HNSW (Malkov and Yashunin, 2018) and IVF-PQ (Jegou et al., 2010; Johnson et al., 2019). This method is efficient for large query ranges. In the extreme case where the query range covers 100% of objects, RF-ANNS becomes equivalent to ANNS, making post-filtering efficient. However, if the query range is small, the ANNS stage may struggle to collect enough qualified candidates, resulting in sub-optimal performance.
Strategy C: Hybrid Filtering. In Strategy A or B, an RF-ANNS query is decomposed into two sub-query systems for attribute filtering and ANNS processing. In contrast, Strategy C employs a single data structure to index and search vectors and attributes simultaneously. Several studies (Gollapudi et al., 2023; Wang et al., 2023) propose to employ state-of-the-art Proximity Graphs (PGs) (Wang et al., 2021a), e.g., HNSW (Malkov and Yashunin, 2018) and Vahama (Jayaram Subramanya et al., 2019; Gollapudi et al., 2023), to implement Strategy C for ANNS with categorical filtration (searching similar vectors whose attributes match a label). The recent study SeRF (Zuo et al., 2024) is the first work implementing Strategy C with PGs for RF-ANNS. It compresses multiple HNSWs into a hybrid index. RF-ANNS is carried out by reconstructing and searching an HNSW index only containing objects in the query range. SeRF achieves state-of-the-art performance on query ranges from 0.3% to 50%. However, it lags behind Strategy A and B for smaller and larger ranges due to the overhead of HNSW reconstruction. Additionally, SeRF lacks support for incremental data insertion, limiting its use in scenarios with new object arrivals.
Problems and Our Solutions. As discussed above, existing methods suffer from two challenges: sub-optimal performance when query range shifts and lack of incremental data insertion support. A trivial solution is to build dedicated indexes for each strategy and adaptively select the best one based on the query range. However, this requires extra effort to ensure data consistency across multiple indexes, leading to high maintenance costs (Wang et al., 2023). To tackle these problems, we propose a UNIFY framework combining Strategies A, B, and C into a unified PG-based index supporting incremental insertion. UNIFY is designed to enhance RF-ANNS performance by prioritizing the following key objectives: (O1) enabling efficient hybrid filtering, (O2) supporting incremental index construction, (O3) integrating pre- and post-filtering strategies, and (O4) implementing a range-aware selection of search strategies. The key techniques in UNIFY to achieve these objectives include:
(1) Segmented Inclusive Graph (SIG) (O1). For efficient hybrid filtering, we introduce a novel graph family named SIG. SIG segments the dataset based on attribute values and theoretically guarantees that the PG of objects from any combination of segments is a sub-graph of SIG. For example, segment a dataset $\mathcal{D}$ into three subsets $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{3}$, and let $\mathbb{G}(\mathcal{X})$ denote the PG constructed on dataset $\mathcal{X}$. The PGs $\mathbb{G}(\mathcal{D}_{1})$, $\mathbb{G}(\mathcal{D}_{2})$, $\mathbb{G}(\mathcal{D}_{1}\cup\mathcal{D}_{2})$, $\dots$, $\mathbb{G}(\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup\mathcal{D}_{3})$ are all included in SIG. This characteristic allows us to reconstruct and search a small PG from relevant segments intersected with the query range to enhance RF-ANNS performance.
(2) Hierarchical Segmented Inclusive Graph (HSIG) (O1 & O2). Based on SIG, we introduce HSIG, a hierarchical graph inspired by HNSW. Similar to HNSW, HSIG is built incrementally. Besides, as a variant of SIG, HSIG can approximately ensure that the HNSW of objects from any combination of segments is a sub-graph of HSIG. By reconstructing and searching an HNSW for relevant segments, we achieve $O(\log(n^{\prime}))$ RF-ANNS time complexity, where $n^{\prime}$ is the number of objects in those segments.
(3) Fusion of skip list connections (O3). The skip list is a classic attribute index optimized for one-dimensional key-value lookup and range search (see Section 2 for more details). We observe that the skip list shares a similar hierarchical structure with our HSIG. Inspired by this, we fuse the skip list connections that reflect the order of attribute values into the hierarchical graph structure to form a hybrid index. By navigating these skip list connections, we can efficiently select objects within the query range, thereby implementing efficient pre-filtering.
(4) Global edge masking (O3). Recall that post-filtering relies on a global ANNS index over the entire dataset. HSIG ensures that the global HNSW is approximately a sub-graph of it, meaning that each node’s global HNSW edges are included within its HSIG edges. We employ an edge masking algorithm to mark the global HNSW edges with a compact bitmap. Navigated by these bitmap-marked edges, we can perform ANNS over the global HNSW to efficiently support post-filtering.
(5) Range-aware search strategy selection (O4). Inspired by the effectiveness of Strategies A, B, and C for different query ranges, we developed a heuristic for range-aware strategy selection. Let $Y$ denote the cardinality of objects that fall within the query range, our heuristic is: use Strategy A for $Y\leq\tau_{A}$, Strategy B for $Y\geq\tau_{B}$, and Strategy C for $\tau_{A}<Y<\tau_{B}$. Here $\tau_{A}$ and $\tau_{B}$ serve as thresholds to distinguish the ranges for which each strategy is most effective, and they can be derived from history data statistics. In this paper, we run a set of sample queries to collect statistics and determine $\tau_{A}$ and $\tau_{B}$. Experimental results show that this heuristic is effective.
Contributions. Our contributions are summarized as follows:
  • We introduced SIG, a novel graph family that segments the dataset based on attribute values, ensuring efficient hybrid filtering by allowing the reconstruction and search of a PG from relevant segments.
  • We developed HSIG, a novel hybrid index that supports efficient hybrid filtering with logarithmic time complexity for RF-ANNS and enables incremental data insertion.
  • We integrated novel auxiliary structures, including skip list connections and edge masking bitmaps, into HSIG to support both pre- and post-filtering strategies. To the best of our knowledge, HSIG is the first index supporting pre-, post-, and hybrid filtering simultaneously.
  • Experiments on real-world datasets demonstrate that our approach significantly outperforms state-of-the-art methods for query ranges from 0.1% to 100% by up to 2.29 times.
2. PRELIMINARIES
2.1. Problem Definition
This paper considers a dataset $\mathcal{D}$ with attributed vectors and nearest neighbors search (NNS) with attribute constraints. Specifically, let $A$ be an attribute (e.g., date, price, or quantity). We use $v[A]$ to denote the attribute value associated with vector $v$. The range filtered nearest neighbors search (RF-NNS) problem is defined as:
Definition 0 (RF-NNS).
Given a dataset $\mathcal{D}$ of $n$ attributed vectors $\{v_{1},v_{2},\dots,v_{n}\}$, a distance function $\Gamma(\cdot,\cdot)$, and a query $Q=(q,[l,h],k)$ with $q$ as the query vector, an integer $k$ from 1 to $n$, and $[l,h]$ a real-valued query range, RF-NNS returns the $kNN(q,\mathcal{R})$, a subset of =$\mathcal{R}=\{v\mid v\in\mathcal{D}$ and $l\leq v[A]\leq h\}$. For any $o\in kNN(q,\mathcal{R})$ and any $u\in\mathcal{R}\setminus kNN(q,\mathcal{R})$, it holds that $\Gamma(o,q)<\Gamma(u,q)$. If $|\mathcal{R}|<k$, all objects in $\mathcal{R}$ are returned.
Due to the ”curse of dimensionality” (Indyk and Motwani, 1998), exact NNS in high-dimensional space is inefficient (Li et al., 2019). As a result, most research focuses on approximate nearest neighbors search (ANNS), which reports approximate results with an optimized recall. Similarly, this paper studies the RF-ANNS problem, which returns an approximate result set $kNN^{\prime}(q,\mathcal{R})$ with an optimized recall:
(1) $recall=\frac{|kNN^{\prime}(\boldsymbol{q},\mathcal{R})\cap kNN(\boldsymbol{q},\mathcal{R})|}{\min(k,|\mathcal{R}|)}.$
Input : $\mathbb{G}$: HNSW layer; $q$: query vector; $ep$: entry point; $k$: an integer.
Output : $q$’s approximate $k$ nearest neighbors on $\mathbb{G}$.
1 push $ep$ to the min-heap $cand$ in the order of distance to $q$;
2 push $ep$ to the max-heap $ann$ in the order of distance to $q$;
3 mark $ep$ as visited;
4 while $|cand|>0$ do
5       $o\leftarrow$ pop the nearest object to $q$ in $cand$;
6       $u\leftarrow$ the furthest object to $q$ in $ann$;
7       if $\Gamma(o,q)>\Gamma(u,q)$ then  break;
8       foreach unvisited $v\in\mathbb{G}[o]$ do
9             mark $v$ as visited;
10             $u\leftarrow$ the furthest object to $q$ in $ann$;
11             if $\Gamma$(v,q) ¡ $\Gamma$(u,q) or $|ann|$ k then
12                   push $v$ to $cand$ and $ann$;
13                   if $|ann|>k$ then pop $ann$;
14
15
16
return $ann$;
Algorithm 1 ANNSearch
2.2. Proximity Graph
A Proximity Graph (PG) (Wang et al., 2021a) treats a vector as a graph node, with connections built based on vector proximity. Various greedy heuristics are proposed to navigate the graph for ANNS (Wang et al., 2021a). In the following, we introduce two PGs related to our work.
$k$ Nearest Neighbor Graph ($k$NNG) (Paredes et al., 2006). Given a dataset $\mathcal{D}$, a $k$NNG is built by connecting each vector $v$ to its $k$ nearest neighbors, $kNN(v,\mathcal{D}\setminus\{v\})$. $k$NNG limits the number of edges per node to at most $k$, making it suitable for memory-constrained environments. However, $k$NNG focuses on local connections and does not guarantee global connectivity, leading to sub-optimal performance compared to state-of-the-art PGs such as HNSW (Malkov and Yashunin, 2018).
Input : $\mathbb{H}$: HNSW; $q$: query vector; $ep$: entry point; $ef$: enlarge factor; $k$: an integer.
Output : $q$’s approximate $k$ nearest neighbors.
1 $L\leftarrow$ max level of $\mathbb{H}$;
2
3foreach $L\geq i\geq 1$ do
       /* Search the $i$-th layer $\mathbb{H}^{i}$. */
4       $ann\leftarrow$ANNSearch($\mathbb{H}^{i}$, $q$, $ep$, 1);
5       $ep\leftarrow$the nearest object to $q$ in $ann$;
6
/* Search the bottom layer $\mathbb{H}^{0}$. */
7 $ann\leftarrow$ANNSearch($\mathbb{H}^{0}$, $q$, $ep$, $ef$);
return top-$k$ nearest objects to $q$ in $ann$;
Algorithm 2 HierarchicalANNS
Input : $\mathbb{G}$: HNSW layer; $v$: the object to insert; $ep$: entry point; $M$: the maximum degree; $efCons$: number of candidate neighbors.
Output : the updated graph.
1 $ann\leftarrow$ ANNSearch($\mathbb{G}$, $v$, $ep$, $efCons$)
2 $\mathbb{G}[v]\leftarrow$ Prune($v$, $ann$, $M$)
3 foreach $o\in\mathbb{G}[v]$ do
4       add $v$ to $\mathbb{G}[o]$;
5       if $|\mathbb{G}[o]|>M$ then  $\mathbb{G}[o]\leftarrow$ Prune($o$, $\mathbb{G}[o]$, $M$) ;
6
return $\mathbb{G}$;
Algorithm 3 InsertLayer
Hierarchical Navigable Small World Graph (HNSW) (Malkov and Yashunin, 2018). HNSW is inspired by the 1D probabilistic structure of the skip list (Pugh, 1990), where each layer is a linked list ordered by 1D values. The bottom layer includes all data objects, while the upper layers contain progressively fewer objects. HNSW extends this structure by replacing linked lists with PGs, enabling efficient hierarchical search. The hierarchical structure of HNSW is similar to that in Figure 4. As outlined in Algorithm 1, HNSW uses a greedy approach to find the $k$NN for query vector $q$ in each layer. The neighbors of each node $v_{i}$ in a given layer $\mathbb{G}$ are stored in adjacency lists $\mathbb{G}[v_{i}]$. The search starts at an entry point $ep$, the most recently inserted vector at the topmost layer (Malkov and Yashunin, 2018). It repeatedly selects the nearest object $o$ to $q$ from $cand$ (Line 5), adds $o$’s unvisited neighbors to $cand$ (Lines 8–12), and updates $ann$ (Lines 11–13). The search terminates when all nodes in $cand$ are farther from $q$ than those in $ann$. As shown in Algorithm 2, the hierarchical search in HNSW begins at a coarse layer to identify promising regions and progressively descends to finer layers for detailed exploration. To improve accuracy, Algorithm 2 uses the parameter $ef$ ($ef>k$), initially exploring $ef$ neighbors to broaden the search region, and finally returns the top-$k$ results.
HNSW is built incrementally based on Algorithm 3. When inserting a new object $v$ into an HNSW layer, the process begins by searching for its top $efCons$ nearest neighbors using the ANNSearch algorithm (Line 1). The Prune heuristic is then applied to limit $v$’s connections to a maximum of $M$ (Line 2). The value of $M$, typically set between 5 and 48 (Malkov and Yashunin, 2018), balances accuracy and efficiency, with its optimal value determined experimentally. The Prune method initially sorts $v$’s candidate neighbors by distance. For each candidate $r$, if there exists a neighbor $e$ that satisfies $\Gamma(v,r)>\Gamma(e,r)$, $r$ is pruned. Otherwise, an edge between $v$ and $r$ is inserted. The process continues until $v$ has $M$ neighbors. Additionally, $v$ is added to the neighbor lists of its identified neighbors, with the Prune method executed for each to maintain the $M$ connection limit (Lines 3–5). The Prune strategy prevents the graph from becoming overly dense, ensuring efficient navigation. HNSW achieves state-of-the-art ANNS time complexity of $O(\log n)$ (Wang et al., 2021a) and demonstrates top-tier practical performance indicated by various benchmarks (Li et al., 2019; Aumüller et al., 2020; Wang et al., 2021a). It also serves as the backbone for various vector databases such as Pinecone (Pinecone, 2021), Weaviate (Weaviate, 2019), and Milvus (Wang et al., 2021b).
3. Segmented Inclusive Graph
We aim to develop a PG-based index that integrates three search strategies and supports incremental construction. We start by discussing the design of a graph optimized for hybrid filtering. In Section 3.1, we introduce the Segmented Inclusive Graph (SIG), a novel graph family that leverages attribute segmentation and graph inclusivity for efficient hybrid filtering. Section 3.2 discusses the challenges of constructing an SIG using the basic PG structure $k$NNG as a case study and presents our practical solutions.
3.1. SIG Overview
Attribute Segmentation. Given a dataset of $n$ objects, RF-ANNS is concerned only with the $n^{\prime}$ objects within the query range, where $n^{\prime}\leq n$. To efficiently filter out the qualified objects, we employ the attribute segmentation method to reduce the search space. Assuming that the attribute distribution remains stable, we first sample objects from the dataset and sort them by their attribute values, then apply an equi-depth histogram (Piatetsky-Shapiro and Connell, 1984) to partition them. The histogram bin boundaries define the attribute intervals for each segment, partitioning the dataset into disjoint subsets with similar size for further indexing.
Graph Inclusivity and SIG. After attribute segmentation, we focus on designing a graph for efficient hybrid filtering. Since a query range intersects a few segments, a straightforward approach is to independently build PGs for each segment. For an RF-ANNS query, we search the local PGs in intersected segments and merge the results for the global $k$NN. However, this method results in search time increasing linearly with the number of intersected segments. State-of-the-art PGs offer an ANNS time complexity of $O(\log n)$, suggesting sub-linear search time growth with the number of intersected segments. For example, with $S$ segments, searching multiple PGs scales as $O(S\log(\frac{n}{S}))$, while searching a single PG containing all objects scales as $O(\log n)$. The ratio between them is $S\left(1-\frac{\log S}{\log n}\right)$. Since $1<S\ll n$, this ratio is greater than 1 but less than $S$, indicating that searching across multiple PGs is less efficient than searching a single PG containing all objects.
Inspired by this, we introduce a novel graph family known as the Segmented Inclusive Graph (SIG). An SIG ensures that the PG of any segment combination is included in (i.e., is a sub-graph of) the SIG. Leveraging this characteristic, termed graph inclusivity, allows efficient hybrid filtering by reconstructing and searching the smaller PG of a few segments covering the query range. Below, we introduce the formal definitions of graph inclusivity and SIG.
Definition 0 (Graph Inclusivity and Segmented Inclusive Graph).
Let $\mathbb{G}(\mathcal{X})$ denote a PG constructed over the dataset $\mathcal{X}$. Given a dataset $\mathcal{D}$ segmented into $S$ disjoint subsets $\mathcal{P}$=$\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{S}$}, a segmented inclusive graph $\mathbb{SIG}(\mathcal{D})$ is a type of graph that possesses the property of graph inclusivity defined as
(2) $\forall r\in\{1,2,\cdots,S\},\ \forall\mathcal{C}\in\mathcal{P}^{(r)},\ \mathbb{G}(\bigcup_{e\in\mathcal{C}}e)\subseteq\mathbb{SIG}(\mathcal{D}),$
where $\mathcal{P}^{(r)}$ denotes all possible combinations of $r$ elements from $\mathcal{P}$.
Example 0.
Assume that $\mathcal{P}=\{\mathcal{D}_{1},\mathcal{D}_{2},\mathcal{D}_{3}\}$, we have $\mathcal{P}^{(2)}=\{\{\mathcal{D}_{1},\mathcal{D}_{2}\},\{\mathcal{D}_{1},\mathcal{D}_{3}\},\{\mathcal{D}_{2},\mathcal{D}_{3}\}\}$. For one of the possible combinations $\mathcal{C}$ $=$ $\{\mathcal{D}_{1},\mathcal{D}_{2}\}$, we have $\bigcup_{e\in\mathcal{C}}(e)$ $=$$\mathcal{D}_{1}\cup\mathcal{D}_{2}$. It holds that $\mathbb{G}(\mathcal{D}_{1}\cup\mathcal{D}_{2})\subseteq\mathbb{SIG}(\mathcal{D})$.
Based on Definition 1, we can derive an exhaustive SIG construction algorithm: build PGs for all possible segment combinations and merge them into a single graph. While this guarantees graph inclusivity, it faces significant computational challenges. With $S$ segments, the number of combinations is $\sum_{i=1}^{S}\binom{S}{i}=2^{S}-1$, and the space complexity of SIG scales as $O(2^{S}-1)$, making the construction of all PGs both computationally and spatially prohibitive [^1] In fact, for RF-ANNS, we only need to consider continuous segments, with potential combinations totaling $0.5(n^{2}-n)$, which is also prohibitive. Since the conclusions in this paper apply to both all segment combinations and continuous ones, we discuss the harder scenario, i.e., all segment combinations, to maintain generality. . Thus, we propose a space-efficient SIG index that scales as $O(S)$ in the next section.
3.2. SIG-$k$NNG
Figure 1. Build multiple $k$NNGs exhaustively $(k=3)$.
Figure 2. Example of the segmented adjacency list ($k=3$).
In this section, we conduct a case study on the basic PG structure $k$NNG to demonstrate the limitations of the exhaustive method. Then, we introduce SIG-$k$NNG, a novel graph structure that guarantees inclusivity without exhaustively building all possible $k$NNGs. Limitation of the Exhaustive Method. For each object $v$ in a dataset $\mathcal{D}$, constructing a $k$NNG involves adding a directed edge $(v,o)$ for each object $o\in kNN(v,\mathcal{D}\setminus\{v\})$. When the dataset is divided into $S$ subsets by attribute segmentation, the exhaustive method requires running the construction algorithm $12^{S}-$ 1 times. Figure 1 illustrates an example. Here, the dataset $\mathcal{D}$ is divided into two subsets $\mathcal{D}_{1}=\{v_{1},v_{2},v_{3},v_{4}\}$ and $\mathcal{D}_{2}=\{v_{5},v_{6},v_{7},v_{8}\}$. We need to build $k$NNGs ($k$=3) for $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{1}\cup\mathcal{D}_{2}$, respectively. Such exhaustive construction is unnecessary, as the $k$NN of a union is inherently within the individual $k$NN sets. For example, all objects in $kNN(v_{4},\mathcal{D}_{1}\cup\mathcal{D}_{2})$ can be found in $kNN(v_{4},\mathcal{D}_{1})$cup $kNN(v_{4},\mathcal{D}_{2})$. In other words, $v_{4}$’s three edges in graph ③ are already included in graphs ① and ②, making the construction of graph ③ redundant.
Structure of SIG-$k$NNG. Inspired by the prior observations, we introduce SIG-$k$NNG, a novel graph structure that achieves inclusivity without the need for exhaustive $k$NNG construction. SIG-$k$NNG uses a segmented adjacency list to storage the outgoing edges for each object in the graph. As illustrated in Figure 2, given a dataset $\mathcal{D}$ segmented into $S$ subsets $\mathcal{D}_{1},\dots,\mathcal{D}_{S}$, the adjacency list for any object $v$ is divided into $S$ chunks. For example, the full adjacency list $\mathbb{G}[v_{4}]$ of object $v_{4}$ is segmented into two chunks $\mathbb{G}[v_{4}][1]$ and $\mathbb{G}[v_{4}][2]$. The $i$-th chunk stores only $v$’s $k$NN within $\mathcal{D}_{i}$. Based on this design, to add SIG-$k$NNG edges for $v$, we perform $k$NN searches $S$ times to find its neighbors in each subset, instead of searching $12^{S}-$ 1 times across all subset combinations. In the following, we introduce the formal definition of SIG-$k$NNG and prove that SIG-$k$NNG can exactly guarantee inclusivity.
Definition 0 (SIG-$k$NNG).
Given a dataset $\mathcal{D}$ segmented into $S$ disjoint subsets $\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{S}$, the SIG-$k$NNG on $\mathcal{D}$ is defined as $\mathbb{F}_{k}(\mathcal{D})=(V_{\mathbb{F}},E_{\mathbb{F}})$, where $V_{\mathbb{F}}=\mathcal{D}$ and $E_{\mathbb{F}}=\bigcup_{v\in\mathcal{D}}\bigcup_{i=1}^{S}E_{i}^{k}(v)$. Here, $E_{i}^{k}(v)$ is the set of edges based on the $i$-th chunk of $v$’s adjacency list, defined as $E_{i}^{k}(v)=\{(v,o)\mid o\in kNN(v,\mathcal{D}_{i}\setminus\{v\})\}$.
Inclusivity of SIG-$k$NNG. We first present the following lemma, which shows that the $k$NN of a union is inherently included in the individual $k$NN sets.
Lemma 0.
Given $r$ disjoint datasets $\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{r}$ and their union set $\mathcal{U}=\bigcup_{i=1}^{r}{\mathcal{D}_{i}}$, it holds that $\forall v\in\mathcal{U}$, $k$NN$(v,\mathcal{U}\setminus\{v\})\subseteq$ $\bigcup_{i=1}^{r}{kNN(v,\mathcal{D}_{i}\setminus\{v\})}$.
Figure 3. Illustration of SIG-$k$NNG’s inlusivity ($k=1$).
Based on Lemma 4, we derive the inclusivity of SIG-$k$NNG, as shown in Theorem 5. The proof is omitted, as it is straightforward to follow.
Theorem 5.
Let $\mathbb{G}_{k}(\mathcal{X})$ and $\mathbb{F}_{k}(\mathcal{X})$ denote a $k$NNG and SIG-$k$NNG for dataset $\mathcal{X}$, respectively. Given a dataset $\mathcal{D}$ segmented into $S$ disjoint subsets $\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{S}$, it follows that SIG-$k$NNG is a segmented inclusive graph. Specifically, for any $r$ distinct integers $i_{1},i_{2},\ldots,i_{r}$ chosen from $[1,S]$ with $1\leq r\leq S$, we have $\mathbb{G}_{k}(\mathcal{U})\subseteq\mathbb{F}_{k}(\mathcal{D})$, where $\mathcal{U}=\bigcup_{j=1}^{r}\mathcal{D}_{i_{j}}$.
Example 0.
Figure 3 demonstrates the inclusivity of SIG-$k$NNG. Given a dataset $\mathcal{D}=\{v_{1},v_{2},\dots,v_{8}\}$, we assume the attribute of $v_{i}$ is $i$ for simplicity. The attribute space $[1,8]$ is divided into two disjoint segments $[1,5)$ and $[5,8]$, leading to two subsets $\mathcal{D}_{1}=\{v_{1},v_{2},v_{3},v_{4}\}$ and $\mathcal{D}_{2}=\{v_{5},v_{6},v_{7},v_{8}\}$. Graphs ①~③ represent the $k$NNGs for $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{1}\cup\mathcal{D}_{2}$, and the SIG-$k$NNG for $\mathcal{D}$ is displayed in graph ④. As illustrated, all three $k$NNGs are sub-graphs of the SIG-$k$NNG.
Figure 4. Illustration of HSIG.
4. Hierarchical Segmented Inclusive Graph
SIG-$k$NNG offers an efficient approach to build an SIG with segmented adjacency lists. Though it theoretically guarantees inclusivity, the underlying $k$NNG is not competitive with state-of-the-art PGs like HNSW (Malkov and Yashunin, 2018). Additionally, SIG-$k$NNG lacks support for incremental insertion. In this section, we expand the basic idea of SIG-$k$NNG, introducing the Hierarchical Segmented Inclusive Graph (HSIG), which uses HNSW as a building block to achieve incremental construction and a logarithmic search complexity.
Figure 4 provides an overview of HSIG, showcasing a hierarchical structure that follows the structure of HNSW and skip lists. HSIG is a unified structure that indexes both vectors and attributes by leveraging the strengths of vector-oriented HNSW and attribute-oriented skip lists. For vector indexing, HSIG organizes the outgoing edges of each object into chunks based on attribute segmentation, inspired by SIG-$k$NNG. Each chunk contains the edges of an HNSW for the corresponding segment. Thus, the backbone graph of HSIG can be seen as a set of HNSWs constructed for different segments (e.g., the three HNSWs with different colors in Figure 4), which are mutually connected with additional edges to ensure inclusivity. Additionally, we incorporate the skip list into the backbone graph to index attributes for efficient pre-filtering and introduce a compact auxiliary structure to optimize post-filtering. The structure and algorithms are detailed in Sections 4.1 and 4.2.
4.1. HSIG Construction
In this section, we describe the core structures of HSIG for hybrid, pre-, and post-filtering sequentially. Then, we integrate them to introduce the complete insertion algorithm.
Backbone Graph for Hybrid Filtering. Based on Theorem 5, we present HSIG’s backbone graph, which is designed to ensure inclusivity regarding HNSW and support efficient hybrid filtering. As depicted in Algorithm 3, when inserting a new object $v$ into an HNSW layer, the core operation involves finding $v$’s $efCons$ nearest neighbors and establishing up to $M$ connections. This procedure is similar to $k$NNG’s construction, which finds $k$NN and establish at most $k$ connections. Therefore, we use the segmented adjacency list introduced by SIG-$k$NNG to guarantee the inclusivity for all possible HNSW connections. With $S$ segments, connections are stored in $S$ chunks, with a maximum degree of $M$ per chunk. The values of $S$ and $M$ are experimentally determined, with default values set to 8 and 16, respectively. Since each chunk contains the HNSW connections in the corresponding segment, the objects and connections in chunk $j$ form a sub-graph $\mathbb{G}_{j}$ of the backbone graph in HSIG. When inserting object $v$ that is located in segment $i$, Algorithm 4 outlines how to build $v$’s connections in segment $j$ at an HSIG layer. First, we search for $v$’s $efCons$ nearest neighbors in $\mathbb{G}_{j}$ (Line 1) using ANNSearch (Algorithm 1), then select up to $M$ neighbors using the Prune algorithm. For each neighbor $o$, we establish mutual connections between $v$ and $o$ within their respective segments (Lines 2–4). If the number of neighbors of $o$ in segment $i$ exceeds $M$, we apply the Prune algorithm to discard the extra neighbors (Lines 5–6). This algorithm guarantees the approximate inclusivity of HSIG, as demonstrated in Section 5.4.
Input : $\mathbb{G}$: HSIG layer; $ep$: entry point; $v$: the object to insert; $i$: index of the segment that $v[A]$ belongs to; $j$: index of the segment to insert; $M$: the maxinum degree; $efCons$: number of candidate neighbors.
Output : $v$’s approximate nearest neighbors in $\mathbb{G}_{j}$ (a sub-graph of $\mathbb{G}$ with only nodes and edges in segment $j$).
1
[…]
3 foreach $o\in$ Prune$(v,ann,M)$ do
4       add $(v,o)$ to $\mathbb{G}[v][j]$;
5       add $(o,v)$ to $\mathbb{G}[o][i]$;
6      if $\mid\mathbb{G}[o][i]\mid>M$ then
7             $\mathbb{G}[o][i]\leftarrow$ Prune($o$, $\mathbb{G}[o][i]$, $M$);
8
9
return $ann$;
Algorithm 4 BackboneConnectionsBuild
Fusing Skip List Connections for Pre-Filtering. For small-range RF-ANNS queries, pre-filtering with an attribute index typically outperforms PG-based methods. This is because PGs prioritize vectors over attributes, making it difficult to filter candidates within the query range. For instance, in the extreme case of a query range containing only one object, using an attribute index to quickly locate the target object is optimal. Thus, we propose fusing an attribute index into the graph structure for effective pre-filtering. As described in Section 2, the skip list is a popular index for efficient 1D key-value lookups and range searches, sharing a hierarchical structure similar to HNSW and HSIG. Inspired by this, we integrate the skip list into the backbone graph of HSIG to form a unified index. Navigated by skip list connections, we can efficiently locate and linearly search objects within the query range. As shown in Figure 4, we use an extra connection $\mathbb{G}[v].next$ to store the ID of $v$’s successor object in the skip list. After inserting an object into the backbone graph via Algorithm 4, we search for its successor in the skip list and store it in $\mathbb{G}[v].next$.
Global Edge Masking for Post-Filtering. For large-range queries, post-filtering with a pure ANNS index is preferable. We use a global HNSW over the entire dataset to support post-filtering. Due to inclusivity, the global HNSW is (approximately) a sub-graph of HSIG. We apply a global edge pruning method to identify these HNSW edges from HSIG edges and use a compact bitmap to mask the unused edges. Since each segment’s sub-graph has a maximum degree of $M$, we similarly limit the connections in the global graph to a maximum of $M$. This procedure is detailed in Algorithm 5. First, we obtain the object $v$’s connections in all segments and select $M$ connections using the Prune algorithm (Line 1). Next, we create a bitmap whose size equals the number of $v$’s neighbors in all segments and set the positions of $v$’s $M$ neighbors to 1 (Line 2). For each neighbor $o$ in the $M$ neighbors (Line 3), if $v$ is also a neighbor of $o$, set the position of $v$ in $o$’s bitmap to 1 (Lines 4–6). Then, we update the bitmap of $o$ (Lines 7–9).
Input : $\mathbb{G}$: HSIG layer; $v$: the object to insert; $M$: the maximum degree.
Output : the updated graph.
1
[…]
3 $\mathbb{G}[v].bitmap\leftarrow$ generate bitmap with $\mathcal{N}$;
4 foreach $o\in\mathcal{N}$ do
5       if $v\in\mathbb{G}[o]$ then
6             $pos\leftarrow$ $v$’s position in $\mathbb{G}[o]$;
7             $\mathbb{G}[o].bitmap[pos]\leftarrow 1$;
8             if  $sum(\mathbb{G}[o].bitmap)>M$ then
9                   $\mathcal{N}^{\prime}\leftarrow$ Prune($o$, $\mathbb{G}[o]$, $M$);
10                   $\mathbb{G}[o].bitmap\leftarrow$ update bitmap with $\mathcal{N}^{\prime}$;
11
12
13
return $\mathbb{G}$;
Algorithm 5 GlobalEdgeMasking
Complete Insertion Algorithm. As aforementioned, the adjacency list of a node in HSIG consists of (1) several chunks for backbone graph connections, (2) a skip list connection, and (3) a bitmap for global connections. The bottom-right of Figure 4 illustrates an example. Here, $\mathbb{G}[v][i]$ is the backbone graph connections of $v$ in the $i$-th segment, $\mathbb{G}[v]$.next stores the ID of $v$’s successor in the skip list, and $\mathbb{G}[v]$.bitmap stores the global edge masks.
Next, we present the complete insertion algorithm to construct an HSIG. The structure of HSIG can be seen as a set of HNSWs for different segments, augmented with auxiliary structures for pre- and post-filtering. The construction consists of three main steps: building the HNSW in each segment, adding skip list connections, and achieving global edge masking. Each step is performed hierarchically and incrementally, as outlined in Algorithm 6. First, we find the segment to which object $v$ belongs (Line 1). The maximum layer $level$ of $v$ is determined randomly using an exponentially decaying probability distribution normalized by $m_{L}$ (Line 2; see (Malkov and Yashunin, 2018) for more details). We traverse all segments (Line 3) to sequentially establish the HNSW $\mathbb{HG}_{j}$ in each segment $j$. We determine the maximum level $L$ and the entry point $ep$ of $\mathbb{HG}_{j}$ (Lines 4–5). For each layer $l$ from $L$ to $level+1$, we use ANNSearch (Algorithm 1) to find the nearest neighbor $ep$ of $v$ in $\mathbb{HG}_{j}^{l}$, where $\mathbb{HG}_{j}^{l}$ denotes layer $l$ in $\mathbb{HG}_{j}$ (Lines 6–7). For each layer $l$ from $level$ to 0, we employ BackboneConnectionsBuild (Algorithm 4) to insert $v$ into $\mathbb{HG}_{j}^{l}$ and use the nearest neighbor of $v$ as the entry point for the next layer (Lines 8–10). After building $\mathbb{HG}_{j}$, we update its entry point if necessary (Line 11) based on the rule that the first object in the topmost layer becomes the entry point. Next, $v$ is inserted into the skip list using the method described in (Pugh, 1990) (Line 12). Finally, we use GlobalEdgeMasking (Algorithm 5) to mask unnecessary edges for post-filtering from layer 0 to $level$ (Lines 13–14).
Input : $\mathbb{HG}$: HSIG; $v$: the object to insert; $S$: number of segments; $M$: the maximum degree; $efCons$: number of candidate neighbors.
Output : the updated HSIG.
1 $i\leftarrow$ ComputeSegmentId($v[A]$);
[…]
3 foreach $1\leq j\leq S$ do
4       $L\leftarrow$ the max level of $\mathbb{HG}_{j}$;
5       $ep\leftarrow$ the entry point of $\mathbb{HG}_{j}$;
6       foreach $L\geq l\geq level+1$ do
7             $ep\leftarrow$ ANNSearch($\mathbb{HG}^{l}_{j}$, $v$, $ep$, 1);
8
9      foreach $level\geq l\geq 0$ do
10             $ann\leftarrow$ BackboneConnectionsBuild($\mathbb{HG}^{l}_{j}$, $ep$, $v$, $i$, $j$, $M$, $efCons$);
11             $ep\leftarrow$ nearest object to $v$ in $ann$;
12
13      update graph entry for $\mathbb{HG}_{j}$ if necessary;
14
15add skip list connections for $v$ in layers $0\dots level$;
16 foreach $0\leq l\leq level$  do
17       $\mathbb{HG}^{l}\leftarrow$ GlobalEdgeMasking($\mathbb{HG}^{l}$, $v$, $M$);
18
return $\mathbb{HG}$;
Algorithm 6 HSIGInsert
4.2. Search on HSIG
To adapt to different query ranges, we propose three search strategies and a range-aware search strategy selection method.
Strategy A (Pre-Filtering). Navigated by the hierarchical skip list connections, we can quickly reach the bottom layer to identify the first object whose attribute value falls within the query range. From there, a linear search is performed to collect the query vector’s $k$NN among those vectors with qualified attribute values. This method is straightforward, so pseudocode is omitted.
Strategy B (Post-Filtering). We employ the hierarchical search scheme outlined in Algorithm 2 to implement ANNS and then filter out objects that fall within the query range. The ANNS is performed over a global HNSW of the entire dataset, with edges marked by bitmaps. During ANNS, we only consider outgoing edges marked as 1 in the object’s bitmap. The search retrieves the top-$ef$ nearest neighbors ($ef>k$) in the bottom layer. We then perform attribute filtering on these top-$ef$ results to obtain the final top-$k$ results.
[…]
Range-aware Strategy Selection. Let $Y$ be the cardinality of objects within a query range. Observing that pre-, post-, and hybrid filtering outperform for small-, large-, and mid-range queries, respectively, we propose the following heuristic for strategy selection: use Strategy A if $Y\leq\tau_{A}$, Strategy B if $Y\geq\tau_{B}$, and Strategy C if $\tau_{A}<Y<\tau_{B}$. Here, $\tau_{A}$ and $\tau_{B}$ are thresholds that distinguish the optimal ranges for each strategy and can be derived from statistical analysis of historical data. Given these thresholds, we can estimate the cardinality of an incoming query to apply the heuristic. Since both statistic collection and cardinality estimation are well studied (Harmouch and Naumann, 2017; Oommen and Rueda, 2002) and are not the focus of this paper, we provide a simple preprocessing method to validate the effectiveness of our heuristic. Specifically, we sample a set of objects from the base dataset to use as queries and assign each a random query range. We then run these queries using the three strategies and record their recall and latency metrics. Given a recall target, we analyze records that meet this requirement, identifying two turning points where pre- and post-filtering outperform hybrid filtering. These two points establish $\tau_{A}$ and $\tau_{B}$, guiding strategy selection for future queries.
Input : $\mathbb{G}$: HSIG layer; $q$: query vector; $[l,h]$: query range; $m$: number of visited neighbors per object; $ep$: entry point; $k$: number of nearest neighbors.
Output : $q$’s approximate $k$ nearest neighbors within $[l,h]$.
1 push $ep$ to the min-heap $cand$ in the order of distance to $q$;
2 push $ep$ to the max-heap $ann$ in the order of distance to $q$;
3 $\mathcal{S}\leftarrow$ segments that intersect with $[l,h]$;
4 while $|cand|>0$ do
5       $o\leftarrow$ pop the nearest object to $q$ in $cand$;
6       $u\leftarrow$ the furthest object to $q$ in $ann$ ;
7       if $\Gamma(o,q)>\Gamma(u,q)$ then  break ;
8       foreach $i\in\mathcal{S}$ do
9             $\mathcal{N}\leftarrow$ the top-($\lceil m/|\mathcal{S}|\rceil$) neighbors in $\mathbb{G}[o][i]$;
10             foreach unvisited $v\in\mathcal{N}$ do
11                   mark $v$ as visited;
12                   $u\leftarrow$ the furthest object to $q$ in $ann$;
13                   if $\Gamma$(v,q) ¡ $\Gamma$(u,q) or $|ann|$ k then
14                         push $v$ to $cand$;
15                         if $l\leq v[A]\leq h$ then  push $v$ to $ann$;
16                         if $|ann|>k$ then pop $ann$ ;
17
18
19
20
return $ann$;
Algorithm 7 HybridFilteringLayer
4.3. Theoretical Analysis
Space Complexity. Following the original work of HNSW (Malkov and Yashunin, 2018), we use a 32-bit integer to store an edge in HSIG and analyze space complexity using 32-bit as a storage unit. In HSIG, each node has up to $MS$ edges (taking $MS$ storage units), a bitmap of size $MS$ (taking $\frac{MS}{32}$ units), and a skip list connection (taking one unit). For a dataset of $n$ objects, the total number of nodes in HSIG is $nL^{\prime}$, where $L^{\prime}$ represents the average number of levels in HSIG. This results in an expected space complexity of $O(nL^{\prime}(\frac{33}{32}MS+1))$. As discussed in (Malkov and Yashunin, 2018), the average number of levels in HNSW is a constant, and since HSIG uses the same strategy to determine the number of levels, $L^{\prime}$ is also a constant in HSIG. The values of $S$ and $M$ are determined experimentally and are generally small constants. Given that $L^{\prime}$, $S$, and $M$ are all considered small constants relative to the dataset size $n$, the space complexity can be simplified to $O(n)$.
Construction Complexity. Consider a dataset with $n$ objects divided into $S$ subsets, each containing $n_{1},n_{2},\dots,n_{S}$ objects. Inserting an object into HSIG requires three operations: (OP1) backbone graph insertion, (OP2) skip list insertion, and (OP3) edge masking. OP3 can be completed in constant time, while OP2 has an expected time complexity of $O(\log n)$ (Pugh, 1990). OP1 involves performing ANNS on the HNSW in each segment, totaling $S$ iterations of ANNS. Since ANNS on an HNSW with $x$ objects takes $O(\log x)$ time, the total cost of OP1 is $\sum_{i=1}^{S}\log(n_{i})$, which is bounded by $\sum_{i=1}^{S}\log n=S\log n$. Therefore, OP1 is an $O(S\log n)$ operation. Neglecting constants, the expected insertion complexity of adding an object is $O(\log n)$, and constructing an HSIG for $n$ objects scales as $O(n\log n)$.
Search Complexity. The complexity scaling of a single search can be strictly analyzed under the assumption that HSIG is able to exactly guarantee inclusivity with respect to HNSW. For an HSIG with $n$ objects, consider a query range covering $Y$ objects and a small constant $k$ that is negligible compared to $n$. The search complexities of the three strategies in HSIG are as follows:
A. Pre-Filtering. Searching the skip list has an expected complexity of $O(\log n)$ (Pugh, 1990), and computing vector distances within the query range takes $O(Y)$ time. Thus, the total time complexity is $O(Y+\log n)$. Due to the range-aware search in HSIG, Pre-Filtering is only employed when $Y$ is smaller than a constant $\tau_{A}$, ensuring that Pre-Filtering typically operates with an $O(\log n)$ complexity.
B. Post-Filtering. This strategy involves searching the global HNSW, which has an expected time complexity of $O(\log n)$ (Malkov and Yashunin, 2018; Wang et al., 2021a).
C. Hybrid Filtering. For RF-ANNS, the hybrid-filtering strategy identifies segments covering the query range and performs ANNS on the HNSW of these segments. Assuming there are $n^{\prime}$ objects in the intersected segments, searching the HNSW takes $O(\log n^{\prime})$ time. Since $n^{\prime}\leq n$, the final expected time complexity is $O(\log n)$. Although the theoretical complexity is derived under the exact inclusivity assumption, experimental results (Section 5.8) confirm the method’s logarithmic scaling with increasing data size, validating its efficiency and scalability for large datasets.
5. EXPERIMENT
5.1. Experimental Setup
Datasets. We use six real-world datasets of varying sizes and dimensions in experiments. The Paper (Wang et al., 2023) and WIT-Image (Zuo et al., 2024) datasets contain both feature vectors and attribute values. The Paper includes publication, topic, and affiliation attributes, which we convert from categorical to numerical, while WIT-Image uses image size as its attribute. For the remaining datasets, which only contain feature vectors, we generate numerical attributes for each object using an attribute generation method similar to that described in (Wang et al., 2021b), augmenting each vector with a randomly generated attribute value between 0 and 10,000. The characteristics of the datasets are detailed in Table 2. For each query, we generate a query range uniformly between 0.1% and 100%.
Table 1. Dataset specifications
Dataset Dimension #Base #Query Type
SIFT1M 128 1,000,000 1,000 Image + Attributes
GIST1M 960 1,000,000 1,000 Image + Attributes
GloVe 100 1,183,514 1,000 Text + Attributes
Msong 420 992,272 200 Audio + Attributes
WIT-Image 2048 1,000,000 1,000 Image + Attributes
Paper 200 2,029,997 10,000 Text + Attributes
Table 2. Index build time and index size.
Method Build Time (s) Index Size (MB)
SIFT1M GIST1M GloVe Msong WIT-Image Paper SIFT1M GIST1M GloVe Msong WIT-Image Paper
Vearch 735 1319 1187 2241 1294 617 692 3905 741 2095 4456 2430
NGT 789 27357 15281 771 8620 814 764 4031 773 1894 4277 2129
NHQ 2806 1689 4956 841 889 5039 78 66 52 99 97 158
ADBV 860 4318 896 2069 10039 2444 21 24 25 22 24 43
Milvus 1459 6931 1560 4289 4983 2477 30 52 35 36 56 61
SeRF 2502 11820 2678 4817 13440 6189 763 3896 704 1852 4185 2096
HSIG 2406 10827 2601 4230 16076 6254 1554 4728 1713 2647 5008 3785
Figure 5. Overall Performance.
Compared Methods. We compare HSIG against six competitors in terms of RF-ANNS performance:
  • ADBV (Wei et al., 2020) is a hybrid analytic engine developed by Alibaba. It enhances PQ (Jegou et al., 2010) for hybrid ANNS and proposes the accuracy-aware, cost-based optimization to generate optimal execution plans.
  • Milvus (Wang et al., 2021b) partitions datasets based on commonly utilized attributes and implements ADBV within each subset.
  • NHQ (Wang et al., 2023) constructs a composite graph index based on the fusion distance of vectors and attributes for hybrid queries. It proposes enhanced edge selection and routing mechanisms to boost query performance.
  • NGT (Yahoo, 2016) is an ANNS library developed by Yahoo Japan that processes hybrid queries using the post-filtering strategy.
  • Vearch (Li et al., 2018; Jingdong, 2020) is a high-dimensional vector retrieval system developed by Jingdong that supports hybrid queries through the post-filtering strategy.
  • SeRF (Zuo et al., 2024) designs a 2D segment graph that compresses multiple ANNS indexes for half-bounded range queries and extends this to support general range queries.
The code for most methods is publicly accessible online. For methods without available code, we implemented them based on their descriptions in the respective papers. Since NHQ initially only supported attribute matching, we modified its fusion distance computation to account for the absolute differences between attribute values and applied the post-filtering strategy for RF-ANNS. We use the Euclidean distance function to measure vector distances.
Metrics. We evaluate query effectiveness by recall and efficiency by measuring the number of queries processed per second (QPS).
Parameter Settings. There are three crucial parameters in HSIG: $S$, $M$, and $efCons$, representing the number of segments, the maximum number of edge connections, and the number of candidate neighbors during index construction, respectively. We use grid search to find the optimal values, setting $S$, $M$, and $efCons$ to 8, 16, and 500, respectively. The parameters $m$ and $ef$ relate to the search process, where $m$ denotes the total number of neighbors visited per object, and $ef$ represents the number of candidates searched during a query. We vary the values of $m$ and $ef$ to generate the recall/QPS curves. We also use grid search to set parameters of baselines.
Implementation Settings. We implement the core parts of the HSIG construction and search algorithms based on hnswlib (Malkov and Yashunin, 2018). The code is written in C++ and compiled with GCC 10.3.1 using the ”-O3” optimization flag. We provide a Python interface for the core indexing library and conduct experiments using Python 3.8.17.
Environment. The scalability experiments are conducted on Alibaba Cloud Linux 3.2104 LTS with 40 physical cores and 512GB memory. The remaining experiments are conducted on a Linux server with the following hardware specifications: an Intel(R) Xeon(R) CPU E5-2609 v3 at 1.90GHz with 6 physical cores, 16GB memory, and the Ubuntu 18.04.5 operating system.
5.2. Overall Performance
We evaluate the query performance of HSIG and its competitors with $k$ values of 10 and 100. Based on the preprocessing method in Section 4.2, we set the default values of $\tau_{A}$ and $\tau_{B}$ to 1% and 50% of the dataset size, respectively. Figure 5 displays the query performance, and Table 2 shows the index sizes and build times. Although HSIG does not excel in index size or build time due to its multi-segment structure and extensive edge connections, it is comparable to SeRF, the state-of-the-art PG-based solution for RF-ANNS. PQ-based methods like ADBV and Milvus are space-efficient but struggle to deliver competitive query accuracy and efficiency compared to PG-based methods. As shown in Figure 5, HSIG consistently outperforms the baselines across all datasets regarding the QPS vs. recall trade-off. For example, with $k=10$ and a recall of around 0.9, HSIG achieves a QPS two orders of magnitude higher than ADBV on the GloVe and GIST1M datasets, one order of magnitude higher than NHQ on the SIFT1M dataset, and outperforms SeRF by up to 2.29 times across all datasets. These results highlight HSIG’s effectiveness due to its unified graph structure and range-aware strategy selection. Additionally, HSIG allows HNSW reconstruction with varying edge degrees using the parameter $m$, whereas SeRF reconstructs HNSW with a fixed edge degree, limiting its query performance. Finally, HSIG outperforms NGT, Vearch, and NHQ, as they employ the post-filtering strategy and perform poorly on small query ranges.
5.3. Effect of Range-aware Search Strategy Selection
In this section, we evaluate HSIG’s performance across small, medium, and large query ranges. The methods compared are as follows: (1) HSIG-pre executes searches on HSIG using the pre-filtering strategy. (2) PQ-pre performs attribute filtering first, followed by PQ-based vector retrieval. (3) Btree-pre uses a B-tree for attribute filtering, followed by brute-force vector retrieval. (4) HSIG-post applies HSIG with the post-filtering strategy. (5) HNSW-post builds HNSW over the entire dataset and performs RF-ANNS using the post-filtering strategy. (6) HSIG-hybrid employs HSIG for RF-ANNS queries using the hybrid filtering strategy. (7) SeRF is a state-of-the-art RF-ANNS solution that uses a hybrid filtering strategy. (8) HSIG-range-aware conducts RF-ANNS using range-aware search strategy selection. (9) Dedicated builds specialized indexes for each strategy, with Btree-pre, HNSW-post, and SeRF used for pre-, post-, and hybrid filtering, respectively, and selects the best strategy based on the query range.
The results are shown in Figure 6. Some methods have missing QPS values for specific query ranges, indicating they could not meet the recall threshold. Notably, HSIG-pre, HSIG-hybrid, and HSIG-post outperform their competitors in small, medium, and large ranges, respectively. For example, HSIG-pre outperforms Btree-pre by 20.3% at a recall of 0.9 in small query ranges. HSIG-hybrid exceeds SeRF by 1.21 times in medium query ranges at a recall of 0.95. HSIG-post outperforms HNSW-post by 37.5% at a recall of 0.99 in large query ranges. Additionally, HSIG-pre outperforms HSIG-hybrid and HSIG-post in small ranges, HSIG-post surpasses HSIG-hybrid and HSIG-pre in large ranges, and HSIG-hybrid performs best among the three strategies in medium ranges, validating the effectiveness of the proposed heuristics for range-aware strategy selection. Finally, HSIG-range-aware consistently outperforms Dedicated by up to 1.1 times across all query ranges, thanks to its unified PG-based index and range-aware strategy selection.
Figure 6. Impact of different query ranges on GloVe dataset.
5.4. Validation of Inclusivity of HSIG
[…]
To further evaluate the impact of inclusivity on query performance, we compare HSIG at varying levels of inclusiveness (30%, 40%, 60%, and 80%) against two competitive methods that guarantee exact inclusivity. The first method, Optimal HNSW, builds an HNSW in real time for objects within each query range. The second method, MS-HNSW, pre-builds HNSWs for each segment. During the search, MS-HNSW identifies the segments intersecting with the query range, retrieves vectors from the corresponding HNSWs, and combines the intermediate results to obtain the final results. The results are shown in Figure 8. While Optimal HNSW offers the best performance, building indexes in real time for every query is time-consuming and impractical. As HSIG’s inclusiveness increases, its query performance improves, approaching that of Optimal HNSW. Additionally, HSIG outperforms MS-HNSW, which requires more distance computations. These results demonstrate the effectiveness of HSIG’s inclusivity, showing that higher inclusiveness leads to better query performance.
Figure 7. Inclusiveness of HSIG.
Figure 8. Impact of Inclusivity.
5.5. Validation of Incremental Insertion
In this section, we evaluate HSIG’s incremental insertion capability. We first build HSIG with 200,000 objects from the SIFT1M dataset, followed by four rounds of incremental insertion, each adding 200,000 objects. We compare the index build time and query performance against the state-of-the-art method SeRF. As shown in Figure 9, HSIG’s build time remains stable with each insertion since the number of inserted objects is consistent, whereas SeRF’s build time increases linearly due to its lack of incremental update support. Moreover, HSIG outperforms SeRF in query efficiency at a recall of 0.95. These results highlight HSIG’s effectiveness in supporting incremental insertions, showing it is suitable for applications with continuously evolving data.
Figure 9. Performance of incremental insertion on SIFT1M.
5.6. Validation of Runtime Neighbor Selection
[…]
Figure 10. Impact of the runtime neighbor selection method.
5.7. Parameter Study
In this section, we analyze the sensitivity of key parameters involved in both index construction and search performance.
Impact of Index Construction Parameters. We analyze the sensitivity of three parameters in HSIG construction: $S$, $M$, and $efCons$. Figures 11 and 12 illustrate how these parameters affect query performance and index build times, respectively. Figure 11a shows the impact of varying $S$ on query performance. We observe a gradual performance improvement as $S$ increases from 2 to 8. However, further increasing $S$ results in a decline in performance because more segments require visiting more neighbors per object during the search. Figure 11b and Figure 11c illustrate the effects of varying $M$ and $efCons$ on query performance, respectively. A lower $M$ degrades the quality of the graph structure, while a higher $M$ increases the number of objects traversed during the search. Similarly, a lower $efCons$ results in insufficient candidate neighbors, whereas a higher $efCons$ includes more irrelevant candidates, both hindering query performance. Therefore, setting these parameters based on the dataset and workload is crucial for balancing query efficiency and accuracy. As shown in Figure 12, index build time increases with higher values of $S$, $M$, and $efCons$ due to the increased distance computations required during construction. Based on a comprehensive evaluation of query performance and index construction time, we select $S=8$, $M=16$, and $efCons=500$ as the default parameter settings.
Figure 11. Performance of index parameters on SIFT1M.
Figure 12. Index build time on different index parameters.
Impact of Search Parameters. Figure 13 shows the impact of varying the parameters $ef$ and $m$ on hybrid filtering in HSIG. Here, $ef$ is typically set to a value greater than $k$. Figure 13a and Figure 13c show that with $m=16$, as $ef$ increases, recall gradually improves while QPS decreases. This occurs because a larger $ef$ requires visiting more objects to gather sufficient candidates, enhancing accuracy but reducing efficiency. Figure 13b and Figure 13d show that with $ef=100$, increasing $m$ leads to a gradual improvement in recall but a decrease in query efficiency. This is because a larger $m$ results in visiting more neighbors of each object during the search, improving accuracy but at the cost of efficiency.
Figure 13. Performance on different search parameters.
Impact of $k$ values. Figure 14 shows the impact of different $k$ values on the hybrid filtering performance of HSIG. HSIG maintains strong query efficiency and accuracy across various $k$ values. However, as $k$ increases from 10 to 100, performance gradually declines due to the increased number of candidates that need to be filtered during the search.
5.8. Scalability
We evaluate the scalability of HSIG using datasets ranging from 10 to 100 million objects. We fix the index parameters to $S=8$, $M=16$, and $efCons=500$, and maintain a query range width of 25%. As shown in Figure 15, both the index size and build time increase almost linearly with the dataset size. Figure 15c plots the hybrid filtering latency versus data size, indicating a logarithmic search complexity. Notably, the recall consistently reaches 0.99 across all dataset sizes. These results demonstrate that HSIG achieves strong scalability in both index construction and query processing.
5.9. Discussions
Range-aware strategy selection. As mentioned in Section 4.2, our range-aware strategy selection method is based on historical data statistics, not query patterns. This approach may face challenges if the data distribution of the base dataset changes significantly over time. To address this, we propose an adaptive method that can detect changes in data distribution. If the change exceeds the user-defined threshold, it resamples objects from the updated base dataset to recalibrate $\tau_{A}$ and $\tau_{B}$, thereby adjusting the range-aware search strategy selection.
RF-ANNS with Multiple Attributes. Existing PG-based indexes struggle to support RF-ANNS queries with multiple attributes, as incorporating multiple attributes into a graph is challenging. However, with two enhancements, HSIG can be extended to handle such queries. While we focus on the case with two attributes in this discussion, the approach can be easily adapted for more attributes. (1) Multiple single-attribute indexes: Build a separate HSIG for each attribute. For a conjunctive query with a query vector $q$, retrieve objects that satisfy $r_{1}(A_{1})$ AND $r_{2}(A_{2})$, where $r_{1}(A_{1})$ and $r_{2}(A_{2})$ are the attribute ranges. The goal is to return the top-$k$ ANN of $q$ among objects satisfying both ranges. Specifically, we modify Line 15 of Algorithm 7 to ”if $v$ satisfies $r_{1}(A_{1})$ AND $r_{2}(A_{2})$, then push $v$ to $ann$”. For disjunctive queries ($r_{1}(A_{1})$ OR $r_{2}(A_{2})$), we use separate indexes for $A_{1}$ and $A_{2}$ and merge the results. (2) Single index for multiple attributes: Create a composite attribute $(A_{1},A_{2})$ and apply the z-order method (Orenstein and Merrett, 1984) to map them into a one-dimensional attribute, enabling the construction of a single HSIG to handle RF-ANNS queries with multiple attributes. We plan to explore a dedicated algorithm for RF-ANNS queries with multiple attributes in future work.
Figure 14. Performance of different $k$ values.
Figure 15. Impact of varying data size on SIFT dataset.
6. RELATED WORK
6.1. ANNS
[…]
6.2. Filtered ANNS
[…]
7. CONCLUSION
This paper addresses RF-ANNS queries over high-dimensional vectors associated with attribute values. Existing methods, including pre-, post-, and hybrid filtering strategies, which apply attribute filtering before, after, or during the ANNS process, suffer performance degradation when query ranges shift. We propose a novel framework called UNIFY, which constructs a unified PG-based index that seamlessly supports all three strategies. Within UNIFY, we introduce SIG, enabling efficient RF-ANNS by reconstructing and searching a PG from relevant segments. Additionally, we present HSIG, a variant of SIG that incorporates a hierarchical structure inspired by HNSW, achieving logarithmic time complexity for RF-ANNS. Experimental results demonstrate that UNIFY outperforms state-of-the-art methods across varying query ranges.

### v2

@@
 PVLDB Reference Format:
-PVLDB, 14(1): XXX-XXX, 2020.
-doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
+PVLDB, 18(4): 1118 - 1130, 2024.
+doi:10.14778/3717755.3717770 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
 Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
@@
 Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
-doi:XX.XX/XXX.XX
+Proceedings of the VLDB Endowment, Vol. 18, No. 4 ISSN 2150-8097.
+doi:10.14778/3717755.3717770
 PVLDB Artifact Availability:
@@
 Let $Y$ denote the cardinality of objects that fall within the query range, our heuristic is: use Strategy A for $Y\leq\tau_{A}$, Strategy B for $Y\geq\tau_{B}$, and Strategy C for $\tau_{A}<Y<\tau_{B}$.
-Here $\tau_{A}$ and $\tau_{B}$ serve as thresholds to distinguish the ranges for which each strategy is most effective, and they can be derived from history data statistics.
+Here, $\tau_{A}$ and $\tau_{B}$ serve as thresholds to distinguish the ranges for which each strategy is most effective, and they can be derived from historical data statistics.
 In this paper, we run a set of sample queries to collect statistics and determine $\tau_{A}$ and $\tau_{B}$.
@@
 HNSW is inspired by the 1D probabilistic structure of the skip list (Pugh, 1990), where each layer is a linked list ordered by 1D values.
-The bottom layer includes all data objects, while the upper layers contain progressively fewer objects.
+The bottom layer includes all objects, while the upper layers contain progressively fewer objects.
 HNSW extends this structure by replacing linked lists with PGs, enabling efficient hierarchical search.
 The hierarchical structure of HNSW is similar to that in Figure 4.
-As outlined in Algorithm 1, HNSW uses a greedy approach to find the $k$NN for query vector $q$ in each layer.
+As shown in Algorithm 1, HNSW uses a greedy approach to find the $k$NN for query vector $q$ in each layer.
 The neighbors of each node $v_{i}$ in a given layer $\mathbb{G}$ are stored in adjacency lists $\mathbb{G}[v_{i}]$.
@@
 It repeatedly selects the nearest object $o$ to $q$ from $cand$ (Line 5), adds $o$’s unvisited neighbors to $cand$ (Lines 8–12), and updates $ann$ (Lines 11–13).
-The search terminates when all nodes in $cand$ are farther from $q$ than those in $ann$.
+The search ends when all nodes in $cand$ are farther from $q$ than those in $ann$.
 As shown in Algorithm 2, the hierarchical search in HNSW begins at a coarse layer to identify promising regions and progressively descends to finer layers for detailed exploration.
@@
 The Prune method initially sorts $v$’s candidate neighbors by distance.
-For each candidate $r$, if there exists a neighbor $e$ that satisfies $\Gamma(v,r)>\Gamma(e,r)$, $r$ is pruned.
+For each candidate $r$, if a neighbor $e$ satisfies $\Gamma(v,r)>\Gamma(e,r)$, $r$ is pruned.
 Otherwise, an edge between $v$ and $r$ is inserted.
 The process continues until $v$ has $M$ neighbors.
-Additionally, $v$ is added to the neighbor lists of its identified neighbors, with the Prune method executed for each to maintain the $M$ connection limit (Lines 3–5).
+Additionally, $v$ is added to the neighbor lists of its identified neighbors, and the Prune method is applied to each neighbor to maintain the $M$-neighbor limit (Lines 3–5).
 The Prune strategy prevents the graph from becoming overly dense, ensuring efficient navigation.
@@
 We aim to develop a PG-based index that integrates three search strategies and supports incremental construction.
-We start by discussing the design of a graph optimized for hybrid filtering.
-In Section 3.1, we introduce the Segmented Inclusive Graph (SIG), a novel graph family that leverages attribute segmentation and graph inclusivity for efficient hybrid filtering.
-Section 3.2 discusses the challenges of constructing an SIG using the basic PG structure $k$NNG as a case study and presents our practical solutions.
+Section 3.1 introduces the Segmented Inclusive Graph (SIG), a novel graph family that leverages attribute segmentation and inclusivity for hybrid filtering.
+Section 3.2 explores the challenges of constructing an SIG using the basic PG structure $k$NNG and presents practical solutions.
 3.1. SIG Overview
@@
 Assuming that the attribute distribution remains stable, we first sample objects from the dataset and sort them by their attribute values, then apply an equi-depth histogram (Piatetsky-Shapiro and Connell, 1984) to partition them.
-The histogram bin boundaries define the attribute intervals for each segment, partitioning the dataset into disjoint subsets with similar size for further indexing.
+The histogram bin boundaries define the attribute intervals for each segment, partitioning the dataset into disjoint subsets with similar sizes for further indexing.
 Graph Inclusivity and SIG.
@@
 It holds that $\mathbb{G}(\mathcal{D}_{1}\cup\mathcal{D}_{2})\subseteq\mathbb{SIG}(\mathcal{D})$.
-Based on Definition 1, we can derive an exhaustive SIG construction algorithm: build PGs for all possible segment combinations and merge them into a single graph.
-While this guarantees graph inclusivity, it faces significant computational challenges.
+Based on Definition 1, we derive an exhaustive SIG construction algorithm: build PGs for all segment combinations and merge them into a single graph.
+While it guarantees graph inclusivity, it faces significant computational challenges.
 With $S$ segments, the number of combinations is $\sum_{i=1}^{S}\binom{S}{i}=2^{S}-1$, and the space complexity of SIG scales as $O(2^{S}-1)$, making the construction of all PGs both computationally and spatially prohibitive [^1] In fact, for RF-ANNS, we only need to consider continuous segments, with potential combinations totaling $0.5(n^{2}-n)$, which is also prohibitive.
@@
 Figure 2. Example of the segmented adjacency list ($k=3$).
-In this section, we conduct a case study on the basic PG structure $k$NNG to demonstrate the limitations of the exhaustive method.
+In this section, we conduct a case study on $k$NNG to demonstrate the limitations of the exhaustive method.
 Then, we introduce SIG-$k$NNG, a novel graph structure that guarantees inclusivity without exhaustively building all possible $k$NNGs.
@@
 Inspired by the prior observations, we introduce SIG-$k$NNG, a novel graph structure that achieves inclusivity without the need for exhaustive $k$NNG construction.
-SIG-$k$NNG uses a segmented adjacency list to storage the outgoing edges for each object in the graph.
+SIG-$k$NNG uses a segmented adjacency list to store the outgoing edges for each object in the graph.
 As illustrated in Figure 2, given a dataset $\mathcal{D}$ segmented into $S$ subsets $\mathcal{D}_{1},\dots,\mathcal{D}_{S}$, the adjacency list for any object $v$ is divided into $S$ chunks.
@@
 Figure 3. Illustration of SIG-$k$NNG’s inlusivity ($k=1$).
-Based on Lemma 4, we derive the inclusivity of SIG-$k$NNG, as shown in Theorem 5. The proof is omitted, as it is straightforward to follow.
+Based on Lemma 4, we derive the inclusivity of SIG-$k$NNG, as presented in Theorem 5. The proof is omitted here for brevity, as it is straightforward to follow.
 Theorem 5.
@@
 As depicted in Algorithm 3, when inserting a new object $v$ into an HNSW layer, the core operation involves finding $v$’s $efCons$ nearest neighbors and establishing up to $M$ connections.
-This procedure is similar to $k$NNG’s construction, which finds $k$NN and establish at most $k$ connections.
+This procedure is similar to $k$NNG’s construction, which finds $k$NN and establishes at most $k$ connections.
 Therefore, we use the segmented adjacency list introduced by SIG-$k$NNG to guarantee the inclusivity for all possible HNSW connections.
@@
 Here, $\mathbb{G}[v][i]$ is the backbone graph connections of $v$ in the $i$-th segment, $\mathbb{G}[v]$.next stores the ID of $v$’s successor in the skip list, and $\mathbb{G}[v]$.bitmap stores the global edge masks.
-Next, we present the complete insertion algorithm to construct an HSIG.
-The structure of HSIG can be seen as a set of HNSWs for different segments, augmented with auxiliary structures for pre- and post-filtering.
-The construction consists of three main steps: building the HNSW in each segment, adding skip list connections, and achieving global edge masking.
+Next, we present the complete insertion algorithm for constructing an HSIG.
+The HSIG structure consists of a set of HNSWs for different segments, augmented with auxiliary structures for pre- and post-filtering.
+Construction involves three main steps: building the HNSW in each segment, adding skip list connections, and applying global edge masking.
 Each step is performed hierarchically and incrementally, as outlined in Algorithm 6.
 First, we find the segment to which object $v$ belongs (Line 1).
-The maximum layer $level$ of $v$ is determined randomly using an exponentially decaying probability distribution normalized by $m_{L}$ (Line 2; see (Malkov and Yashunin, 2018) for more details).
-We traverse all segments (Line 3) to sequentially establish the HNSW $\mathbb{HG}_{j}$ in each segment $j$.
-We determine the maximum level $L$ and the entry point $ep$ of $\mathbb{HG}_{j}$ (Lines 4–5).
+The maximum layer $level$ of $v$ is randomly assigned using an exponentially decaying probability distribution normalized by $m_{L}$ (Line 2; see (Malkov and Yashunin, 2018) for details).
+We then traverse all segments (Line 3) to sequentially build the HNSW $\mathbb{HG}_{j}$ in each segment $j$, determining its maximum level $L$ and entry point $ep$ (Lines 4–5).
 For each layer $l$ from $L$ to $level+1$, we use ANNSearch (Algorithm 1) to find the nearest neighbor $ep$ of $v$ in $\mathbb{HG}_{j}^{l}$, where $\mathbb{HG}_{j}^{l}$ denotes layer $l$ in $\mathbb{HG}_{j}$ (Lines 6–7).
-For each layer $l$ from $level$ to 0, we employ BackboneConnectionsBuild (Algorithm 4) to insert $v$ into $\mathbb{HG}_{j}^{l}$ and use the nearest neighbor of $v$ as the entry point for the next layer (Lines 8–10).
-After building $\mathbb{HG}_{j}$, we update its entry point if necessary (Line 11) based on the rule that the first object in the topmost layer becomes the entry point.
-Next, $v$ is inserted into the skip list using the method described in (Pugh, 1990) (Line 12).
+Then, for each layer $l$ from $level$ to 0, we employ BackboneConnectionsBuild (Algorithm 4) to insert $v$ into $\mathbb{HG}_{j}^{l}$, using its nearest neighbor as the entry point for the next layer (Lines 8–10).
+After building $\mathbb{HG}_{j}$, we update its entry point if necessary (Line 11), ensuring the first object in the topmost layer serves as the entry.
+Next, $v$ is inserted into the skip list using the method in (Pugh, 1990) (Line 12).
 Finally, we use GlobalEdgeMasking (Algorithm 5) to mask unnecessary edges for post-filtering from layer 0 to $level$ (Lines 13–14).
@@
 Strategy B (Post-Filtering).
-We employ the hierarchical search scheme outlined in Algorithm 2 to implement ANNS and then filter out objects that fall within the query range.
-The ANNS is performed over a global HNSW of the entire dataset, with edges marked by bitmaps.
-During ANNS, we only consider outgoing edges marked as 1 in the object’s bitmap.
-The search retrieves the top-$ef$ nearest neighbors ($ef>k$) in the bottom layer.
-We then perform attribute filtering on these top-$ef$ results to obtain the final top-$k$ results.
+We employ the hierarchical search scheme in Algorithm 2 for ANNS, followed by filtering objects within the query range.
+ANNS is performed over a global HNSW with edges marked by bitmaps.
+During the search, only outgoing edges marked as 1 in an object’s bitmap are considered.
+The search retrieves the top-$ef$ nearest neighbors ($ef>k$) in the bottom layer, after which attribute filtering is applied to obtain the top-$k$ results.
 […]
@@
 Let $Y$ be the cardinality of objects within a query range.
-Observing that pre-, post-, and hybrid filtering outperform for small-, large-, and mid-range queries, respectively, we propose the following heuristic for strategy selection: use Strategy A if $Y\leq\tau_{A}$, Strategy B if $Y\geq\tau_{B}$, and Strategy C if $\tau_{A}<Y<\tau_{B}$.
-Here, $\tau_{A}$ and $\tau_{B}$ are thresholds that distinguish the optimal ranges for each strategy and can be derived from statistical analysis of historical data.
-Given these thresholds, we can estimate the cardinality of an incoming query to apply the heuristic.
-Since both statistic collection and cardinality estimation are well studied (Harmouch and Naumann, 2017; Oommen and Rueda, 2002) and are not the focus of this paper, we provide a simple preprocessing method to validate the effectiveness of our heuristic.
-Specifically, we sample a set of objects from the base dataset to use as queries and assign each a random query range.
-We then run these queries using the three strategies and record their recall and latency metrics.
-Given a recall target, we analyze records that meet this requirement, identifying two turning points where pre- and post-filtering outperform hybrid filtering.
-These two points establish $\tau_{A}$ and $\tau_{B}$, guiding strategy selection for future queries.
+Observing that pre-, post-, and hybrid filtering perform best for small-, large-, and mid-range queries, respectively, we propose the following heuristic for strategy selection: use Strategy A if $Y\leq\tau_{A}$, Strategy B if $Y\geq\tau_{B}$, and Strategy C if $\tau_{A}<Y<\tau_{B}$.
+Here, $\tau_{A}$ and $\tau_{B}$ are thresholds distinguishing the optimal ranges for each strategy, derived from historical data analysis.
+Given these thresholds, we estimate the cardinality of an incoming query to apply the heuristic.
+Since statistic collection and cardinality estimation are well studied (Harmouch and Naumann, 2017; Oommen and Rueda, 2002) and not the focus of this paper, we provide a simple preprocessing method to validate our heuristic.
+Specifically, we sample objects from the base dataset as queries and assign each a random query range.
+We then execute these queries using the three strategies and record recall and latency metrics.
+Given a recall target, we analyze records meeting this requirement and identify two turning points where pre- and post-filtering outperform hybrid filtering.
+These points establish $\tau_{A}$ and $\tau_{B}$, guiding strategy selection for future queries.
 Input : $\mathbb{G}$: HSIG layer; $q$: query vector; $[l,h]$: query range; $m$: number of visited neighbors per object; $ep$: entry point; $k$: number of nearest neighbors.
@@
 Space Complexity.
-Following the original work of HNSW (Malkov and Yashunin, 2018), we use a 32-bit integer to store an edge in HSIG and analyze space complexity using 32-bit as a storage unit.
-In HSIG, each node has up to $MS$ edges (taking $MS$ storage units), a bitmap of size $MS$ (taking $\frac{MS}{32}$ units), and a skip list connection (taking one unit).
-For a dataset of $n$ objects, the total number of nodes in HSIG is $nL^{\prime}$, where $L^{\prime}$ represents the average number of levels in HSIG.
+Following HNSW (Malkov and Yashunin, 2018), we use a 32-bit integer to store each edge in HSIG and analyze space complexity using 32-bit as a storage unit.
+In HSIG, each node has up to $MS$ edges (taking $MS$ units), a bitmap of size $MS$ (taking $\frac{MS}{32}$ units), and a skip list connection (taking one unit).
+For a dataset of $n$ objects, HSIG contains $nL^{\prime}$ nodes, where $L^{\prime}$ is the average number of levels.
 This results in an expected space complexity of $O(nL^{\prime}(\frac{33}{32}MS+1))$.
 As discussed in (Malkov and Yashunin, 2018), the average number of levels in HNSW is a constant, and since HSIG uses the same strategy to determine the number of levels, $L^{\prime}$ is also a constant in HSIG.
-The values of $S$ and $M$ are determined experimentally and are generally small constants.
-Given that $L^{\prime}$, $S$, and $M$ are all considered small constants relative to the dataset size $n$, the space complexity can be simplified to $O(n)$.
+$S$ and $M$ are experimentally determined and are generally small constants.
+Given that $L^{\prime}$, $S$, and $M$ are all considered small constants relative to $n$, the space complexity can be simplified to $O(n)$.
 Construction Complexity.
@@
 Inserting an object into HSIG requires three operations: (OP1) backbone graph insertion, (OP2) skip list insertion, and (OP3) edge masking.
-OP3 can be completed in constant time, while OP2 has an expected time complexity of $O(\log n)$ (Pugh, 1990).
-OP1 involves performing ANNS on the HNSW in each segment, totaling $S$ iterations of ANNS.
-Since ANNS on an HNSW with $x$ objects takes $O(\log x)$ time, the total cost of OP1 is $\sum_{i=1}^{S}\log(n_{i})$, which is bounded by $\sum_{i=1}^{S}\log n=S\log n$.
-Therefore, OP1 is an $O(S\log n)$ operation.
-Neglecting constants, the expected insertion complexity of adding an object is $O(\log n)$, and constructing an HSIG for $n$ objects scales as $O(n\log n)$.
+OP3 runs in constant time, while OP2 has an expected complexity of $O(\log n)$ (Pugh, 1990).
+OP1 performs ANNS on the HNSW of each segment, requiring $S$ ANNS iterations.
+Since ANNS on an HNSW with $x$ objects takes $O(\log x)$ time, OP1 costs $\sum_{i=1}^{S}\log(n_{i})$ bounded by $S\log n$, making OP1 an $O(S\log n)$ operation.
+Neglecting constants, the expected complexity of inserting an object is $O(\log n)$, and constructing an HSIG for $n$ objects scales as $O(n\log n)$.
 Search Complexity.
-The complexity scaling of a single search can be strictly analyzed under the assumption that HSIG is able to exactly guarantee inclusivity with respect to HNSW.
-For an HSIG with $n$ objects, consider a query range covering $Y$ objects and a small constant $k$ that is negligible compared to $n$.
+The complexity scaling of a single search can be strictly analyzed under the assumption that HSIG exactly guarantees inclusivity with respect to HNSW.
+For an HSIG with $n$ objects, consider a query range covering $Y$ objects and a small constant $k$ negligible compared to $n$.
 The search complexities of the three strategies in HSIG are as follows:
@@
 Searching the skip list has an expected complexity of $O(\log n)$ (Pugh, 1990), and computing vector distances within the query range takes $O(Y)$ time.
-Thus, the total time complexity is $O(Y+\log n)$.
-Due to the range-aware search in HSIG, Pre-Filtering is only employed when $Y$ is smaller than a constant $\tau_{A}$, ensuring that Pre-Filtering typically operates with an $O(\log n)$ complexity.
+Thus, the total complexity is $O(Y+\log n)$.
+Due to range-aware search in HSIG, Pre-Filtering is used only when $Y$ is smaller than a constant $\tau_{A}$, ensuring it typically operates with an $O(\log n)$ complexity.
 B. Post-Filtering. This strategy involves searching the global HNSW, which has an expected time complexity of $O(\log n)$ (Malkov and Yashunin, 2018; Wang et al., 2021a).
@@
 Hybrid Filtering.
-For RF-ANNS, the hybrid-filtering strategy identifies segments covering the query range and performs ANNS on the HNSW of these segments.
-Assuming there are $n^{\prime}$ objects in the intersected segments, searching the HNSW takes $O(\log n^{\prime})$ time.
-Since $n^{\prime}\leq n$, the final expected time complexity is $O(\log n)$.
-Although the theoretical complexity is derived under the exact inclusivity assumption, experimental results (Section 5.8) confirm the method’s logarithmic scaling with increasing data size, validating its efficiency and scalability for large datasets.
-5. EXPERIMENT
-5.1. Experimental Setup
-Datasets.
-We use six real-world datasets of varying sizes and dimensions in experiments.
-The Paper (Wang et al., 2023) and WIT-Image (Zuo et al., 2024) datasets contain both feature vectors and attribute values.
-The Paper includes publication, topic, and affiliation attributes, which we convert from categorical to numerical, while WIT-Image uses image size as its attribute.
-For the remaining datasets, which only contain feature vectors, we generate numerical attributes for each object using an attribute generation method similar to that described in (Wang et al., 2021b), augmenting each vector with a randomly generated attribute value between 0 and 10,000.
-The characteristics of the datasets are detailed in Table 2.
-For each query, we generate a query range uniformly between 0.1% and 100%.
+For RF-ANNS, hybrid filtering selects segments covering the query range and performs ANNS on their HNSWs.
+Let $n^{\prime}$ be the number of objects in the intersected segments, searching the HNSW takes $O(\log n^{\prime})$ time.
+Since $n^{\prime}\leq n$, the expected time complexity is $O(\log n)$.
+Although the theoretical complexity is derived under the exact inclusivity assumption, experimental results (Section 5.8) confirm the method’s logarithmic scaling with data size, validating its efficiency and scalability.
 Table 1. Dataset specifications
@@
 Paper 200 2,029,997 10,000 Text + Attributes
+5. EXPERIMENT
+5.1. Experimental Setup
+Datasets.
+We use six real-world datasets of varying sizes and dimensions.
+The Paper (Wang et al., 2023) and WIT-Image (Zuo et al., 2024) datasets include both feature vectors and attributes.
+The Paper contains publication, topic, and affiliation attributes, which we convert from categorical to numerical, while WIT-Image uses image size as its attribute.
+For the remaining datasets, which contain only feature vectors, we generate numerical attributes using a method similar to (Wang et al., 2021b), assigning each vector a random value between 0 and 10,000.
+The dataset characteristics are detailed in Table 1.
+For each query, we generate a query range uniformly between 0.1% and 100%.
 Table 2. Index build time and index size.
@@
   • SeRF (Zuo et al., 2024) designs a 2D segment graph that compresses multiple ANNS indexes for half-bounded range queries and extends this to support general range queries.
-The code for most methods is publicly accessible online.
-For methods without available code, we implemented them based on their descriptions in the respective papers.
-Since NHQ initially only supported attribute matching, we modified its fusion distance computation to account for the absolute differences between attribute values and applied the post-filtering strategy for RF-ANNS.
 We use the Euclidean distance function to measure vector distances.
@@
 Parameter Settings.
-There are three crucial parameters in HSIG: $S$, $M$, and $efCons$, representing the number of segments, the maximum number of edge connections, and the number of candidate neighbors during index construction, respectively.
-We use grid search to find the optimal values, setting $S$, $M$, and $efCons$ to 8, 16, and 500, respectively.
-The parameters $m$ and $ef$ relate to the search process, where $m$ denotes the total number of neighbors visited per object, and $ef$ represents the number of candidates searched during a query.
-We vary the values of $m$ and $ef$ to generate the recall/QPS curves.
-We also use grid search to set parameters of baselines.
+The parameters $S$, $M$, and $efCons$ represent the number of segments, maximum edge connections, and candidate neighbors during index construction, respectively.
+We use grid search to determine their optimal values, setting $S$, $M$, and $efCons$ to 8, 16, and 500, respectively.
+The parameters $m$ and $ef$ relate to the search process, where $m$ is the total number of neighbors visited per object, and $ef$ is the number of candidates searched during a query.
+We vary $m$ and $ef$ to generate recall/QPS curves and apply grid search to set baseline parameters.
 Implementation Settings.
-We implement the core parts of the HSIG construction and search algorithms based on hnswlib (Malkov and Yashunin, 2018).
+We implement HSIG construction and search algorithms based on hnswlib (Malkov and Yashunin, 2018).
 The code is written in C++ and compiled with GCC 10.3.1 using the ”-O3” optimization flag.
-We provide a Python interface for the core indexing library and conduct experiments using Python 3.8.17.
+A Python interface is provided for the indexing library, and experiments are conducted using Python 3.8.17.
 Environment.
-The scalability experiments are conducted on Alibaba Cloud Linux 3.2104 LTS with 40 physical cores and 512GB memory.
-The remaining experiments are conducted on a Linux server with the following hardware specifications: an Intel(R) Xeon(R) CPU E5-2609 v3 at 1.90GHz with 6 physical cores, 16GB memory, and the Ubuntu 18.04.5 operating system.
+Scalability experiments are conducted on Alibaba Cloud Linux 3.2104 LTS with 40 cores and 512GB memory.
+Other experiments are conducted on a Linux server with an Intel(R) Xeon(R) E5-2609 v3 (1.90GHz, 6 cores), 16GB memory, and Ubuntu 18.04.5.
 5.2. Overall Performance
 We evaluate the query performance of HSIG and its competitors with $k$ values of 10 and 100.
-Based on the preprocessing method in Section 4.2, we set the default values of $\tau_{A}$ and $\tau_{B}$ to 1% and 50% of the dataset size, respectively.
-Figure 5 displays the query performance, and Table 2 shows the index sizes and build times.
+Based on the preprocessing method in Section 4.2, we set $\tau_{A}$ and $\tau_{B}$ to 1% and 50% of the dataset size.
+Figure 5 shows query performance, while Table 2 presents index sizes and build times.
 Although HSIG does not excel in index size or build time due to its multi-segment structure and extensive edge connections, it is comparable to SeRF, the state-of-the-art PG-based solution for RF-ANNS.
-PQ-based methods like ADBV and Milvus are space-efficient but struggle to deliver competitive query accuracy and efficiency compared to PG-based methods.
-As shown in Figure 5, HSIG consistently outperforms the baselines across all datasets regarding the QPS vs.
+PQ-based methods like ADBV and Milvus are space-efficient but struggle with query accuracy and efficiency compared to PG-based methods.
+As shown in Figure 5, HSIG consistently outperforms baselines across all datasets regarding the QPS vs.
 recall trade-off.
 For example, with $k=10$ and a recall of around 0.9, HSIG achieves a QPS two orders of magnitude higher than ADBV on the GloVe and GIST1M datasets, one order of magnitude higher than NHQ on the SIFT1M dataset, and outperforms SeRF by up to 2.29 times across all datasets.
-These results highlight HSIG’s effectiveness due to its unified graph structure and range-aware strategy selection.
+These results highlight HSIG’s effectiveness, benefiting from its unified graph structure and range-aware strategy selection.
 Additionally, HSIG allows HNSW reconstruction with varying edge degrees using the parameter $m$, whereas SeRF reconstructs HNSW with a fixed edge degree, limiting its query performance.
@@
 5.3. Effect of Range-aware Search Strategy Selection
-In this section, we evaluate HSIG’s performance across small, medium, and large query ranges.
-The methods compared are as follows: (1) HSIG-pre executes searches on HSIG using the pre-filtering strategy.
+We evaluate HSIG’s performance across small, medium, and large query ranges.
+The methods compared are as follows: (1) HSIG-pre uses HSIG with the pre-filtering strategy.
 (2) PQ-pre performs attribute filtering first, followed by PQ-based vector retrieval.
@@
 (4) HSIG-post applies HSIG with the post-filtering strategy.
-(5) HNSW-post builds HNSW over the entire dataset and performs RF-ANNS using the post-filtering strategy.
+(5) HNSW-post builds HNSW over the entire dataset and applies the post-filtering strategy.
 (6) HSIG-hybrid employs HSIG for RF-ANNS queries using the hybrid filtering strategy.
-(7) SeRF is a state-of-the-art RF-ANNS solution that uses a hybrid filtering strategy.
-(8) HSIG-range-aware conducts RF-ANNS using range-aware search strategy selection.
-(9) Dedicated builds specialized indexes for each strategy, with Btree-pre, HNSW-post, and SeRF used for pre-, post-, and hybrid filtering, respectively, and selects the best strategy based on the query range.
+(7) SeRF is a state-of-the-art RF-ANNS solution that uses a hybrid filtering.
+(8) HSIG-range-aware uses HSIG with range-aware search strategy selection.
+(9) Dedicated builds specialized indexes for each strategy, with Btree-pre, HNSW-post, and SeRF used for pre-, post-, and hybrid filtering, respectively, and selects the best strategy based on query range.
 The results are shown in Figure 6.
 Some methods have missing QPS values for specific query ranges, indicating they could not meet the recall threshold.
-Notably, HSIG-pre, HSIG-hybrid, and HSIG-post outperform their competitors in small, medium, and large ranges, respectively.
+HSIG-pre, HSIG-hybrid, and HSIG-post outperform competitors in small, medium, and large ranges, respectively.
 For example, HSIG-pre outperforms Btree-pre by 20.3% at a recall of 0.9 in small query ranges.
 HSIG-hybrid exceeds SeRF by 1.21 times in medium query ranges at a recall of 0.95.
-HSIG-post outperforms HNSW-post by 37.5% at a recall of 0.99 in large query ranges.
-Additionally, HSIG-pre outperforms HSIG-hybrid and HSIG-post in small ranges, HSIG-post surpasses HSIG-hybrid and HSIG-pre in large ranges, and HSIG-hybrid performs best among the three strategies in medium ranges, validating the effectiveness of the proposed heuristics for range-aware strategy selection.
+HSIG-post surpasses HNSW-post by 37.5% at a recall of 0.99 in large query ranges.
+Additionally, HSIG-pre outperforms HSIG-hybrid and HSIG-post in small ranges, HSIG-post surpasses HSIG-hybrid and HSIG-pre in large ranges, and HSIG-hybrid performs best among the three strategies in medium ranges, validating our range-aware heuristics.
 Finally, HSIG-range-aware consistently outperforms Dedicated by up to 1.1 times across all query ranges, thanks to its unified PG-based index and range-aware strategy selection.
@@
 As HSIG’s inclusiveness increases, its query performance improves, approaching that of Optimal HNSW.
-Additionally, HSIG outperforms MS-HNSW, which requires more distance computations.
+Additionally, HSIG consistently outperforms MS-HNSW, which requires more distance computations.
 These results demonstrate the effectiveness of HSIG’s inclusivity, showing that higher inclusiveness leads to better query performance.
@@
 Figure 8. Impact of Inclusivity.
+Figure 9. Performance of incremental insertion on SIFT1M.
 5.5. Validation of Incremental Insertion
@@
 Moreover, HSIG outperforms SeRF in query efficiency at a recall of 0.95.
-These results highlight HSIG’s effectiveness in supporting incremental insertions, showing it is suitable for applications with continuously evolving data.
-Figure 9. Performance of incremental insertion on SIFT1M.
+These results highlight HSIG’s effectiveness in supporting incremental insertions, showing that it is suitable for applications with continuously evolving data.
 5.6. Validation of Runtime Neighbor Selection
@@
 Figure 10. Impact of the runtime neighbor selection method.
+Figure 11. Performance of index parameters on SIFT1M.
 5.7. Parameter Study
-In this section, we analyze the sensitivity of key parameters involved in both index construction and search performance.
 Impact of Index Construction Parameters.
 We analyze the sensitivity of three parameters in HSIG construction: $S$, $M$, and $efCons$.
-Figures 11 and 12 illustrate how these parameters affect query performance and index build times, respectively.
-Figure 11a shows the impact of varying $S$ on query performance.
-We observe a gradual performance improvement as $S$ increases from 2 to 8.
-However, further increasing $S$ results in a decline in performance because more segments require visiting more neighbors per object during the search.
-Figure 11b and Figure 11c illustrate the effects of varying $M$ and $efCons$ on query performance, respectively.
-A lower $M$ degrades the quality of the graph structure, while a higher $M$ increases the number of objects traversed during the search.
-Similarly, a lower $efCons$ results in insufficient candidate neighbors, whereas a higher $efCons$ includes more irrelevant candidates, both hindering query performance.
+Figures 11 and 12 show how parameters affect query performance and index build times, respectively.
+Figure 11a shows the impact of varying $S$ on query performance, with a performance increase from 2 to 8, followed by a decline as $S$ grows further due to increased neighbor visits.
+Figure 11b and Figure 11c illustrate the effects of varying $M$ and $efCons$ on query performance.
+Lower $M$ degrades graph quality, while higher $M$ increases the number of objects traversed during the search.
+Similarly, lower $efCons$ leads to insufficient candidates, whereas higher $efCons$ includes irrelevant candidates, both hindering query performance.
 Therefore, setting these parameters based on the dataset and workload is crucial for balancing query efficiency and accuracy.
-As shown in Figure 12, index build time increases with higher values of $S$, $M$, and $efCons$ due to the increased distance computations required during construction.
-Based on a comprehensive evaluation of query performance and index construction time, we select $S=8$, $M=16$, and $efCons=500$ as the default parameter settings.
-Figure 11. Performance of index parameters on SIFT1M.
+As shown in Figure 12, index build time rises with higher $S$, $M$, and $efCons$ due to increased distance computations.
+Based on a comprehensive evaluation of query performance and index construction time, we select $S=8$, $M=16$, and $efCons=500$ as default settings.
 Figure 12. Index build time on different index parameters.
@@
 Impact of $k$ values.
-Figure 14 shows the impact of different $k$ values on the hybrid filtering performance of HSIG.
-HSIG maintains strong query efficiency and accuracy across various $k$ values.
+Figure 14 shows the impact of different $k$ values on query performance of HSIG.
+HSIG maintains strong efficiency and accuracy across various $k$ values.
 However, as $k$ increases from 10 to 100, performance gradually declines due to the increased number of candidates that need to be filtered during the search.
@@
 This approach may face challenges if the data distribution of the base dataset changes significantly over time.
-To address this, we propose an adaptive method that can detect changes in data distribution.
-If the change exceeds the user-defined threshold, it resamples objects from the updated base dataset to recalibrate $\tau_{A}$ and $\tau_{B}$, thereby adjusting the range-aware search strategy selection.
+To address this issue, we propose an adaptive method that can detect changes in data distribution.
+If the change exceeds the user-defined threshold, the method resamples objects from the updated base dataset to recalibrate $\tau_{A}$ and $\tau_{B}$, thereby adjusting the range-aware search strategy selection.
 RF-ANNS with Multiple Attributes.
@@
 However, with two enhancements, HSIG can be extended to handle such queries.
-While we focus on the case with two attributes in this discussion, the approach can be easily adapted for more attributes.
+While we use the case with two attributes in this discussion, the method can be adapted for more attributes.
 (1) Multiple single-attribute indexes: Build a separate HSIG for each attribute.
 For a conjunctive query with a query vector $q$, retrieve objects that satisfy $r_{1}(A_{1})$ AND $r_{2}(A_{2})$, where $r_{1}(A_{1})$ and $r_{2}(A_{2})$ are the attribute ranges.
-The goal is to return the top-$k$ ANN of $q$ among objects satisfying both ranges.
+We aim to return the top-$k$ ANN of $q$ among objects satisfying both ranges.
 Specifically, we modify Line 15 of Algorithm 7 to ”if $v$ satisfies $r_{1}(A_{1})$ AND $r_{2}(A_{2})$, then push $v$ to $ann$”.
@@
 Figure 14. Performance of different $k$ values.
-Figure 15. Impact of varying data size on SIFT dataset.
 6. RELATED WORK
@@
 […]
+Figure 15. Impact of varying data size on SIFT dataset.
 6.2. Filtered ANNS
@@
 Experimental results demonstrate that UNIFY outperforms state-of-the-art methods across varying query ranges.
+Acknowledgements.
+Bin Yao was supported by the Oceanic Interdisciplinary Program of Shanghai Jiao Tong University (SL2023ZD102) and Alibaba Group through Alibaba Innovative Research (AIR) Program.
+Zhongpu Chen was supported by Sichuan Science and Technology Program (2024NSFSC1460).